### v1

# Optimizing Probabilistic Conformal Prediction with Vectorized Non-Conformity Scores

Abstract
Generative models have shown significant promise in critical domains such as medical diagnosis, autonomous driving, and climate science, where reliable decision-making hinges on accurate uncertainty quantification. While probabilistic conformal prediction (PCP) offers a powerful framework for this purpose, its coverage efficiency – the size of the uncertainty set – is limited when dealing with complex underlying distributions and a finite number of generated samples. In this paper, we propose a novel PCP framework that enhances efficiency by first vectorizing the non-conformity scores with ranked samples and then optimizing the shape of the prediction set by varying the quantiles for samples at the same rank. Our method delivers valid coverage while producing discontinuous and more efficient prediction sets, making it particularly suited for high-stakes applications. We demonstrate the effectiveness of our approach through experiments on both synthetic and real-world datasets.
1 Introduction
Uncertainty quantification is a critical component in predictive modeling and decision-making processes across a wide range of fields, including finance (Fujimoto et al.,, 2022; Blasco et al.,, 2024), healthcare (Seoni et al.,, 2023), and engineering (Kumar et al.,, 2023). In real-world applications, models often encounter inherent uncertainties due to data noise, model limitations, or unforeseen events. Accurately quantifying these uncertainties enables practitioners to assess risks and make informed decisions. For instance, in medical diagnostics, providing uncertainty estimates alongside disease predictions can guide clinicians in determining the necessity of further tests or interventions (Zou et al.,, 2023; Lambert et al.,, 2024). Similarly, in autonomous driving, understanding the uncertainty in object detection can enhance safety by prompting caution in ambiguous situations Michelmore et al., (2020); Su et al., (2023); Grewal et al., (2024).
Conformal prediction (Vovk et al.,, 2005) has emerged as a powerful framework for uncertainty quantification, offering statistically valid prediction sets without making distributional assumptions. Its popularity stems from its ability to provide finite-sample guarantees under the assumption of data exchangeability, making it widely applicable in various machine learning tasks. For example, in regression analysis, conformal prediction can generate prediction intervals that adapt to the local variability of the data, thereby offering more reliable uncertainty estimates compared to traditional methods.
Recent advancements, such as probabilistic conformal prediction (PCP) using conditional random samples (Wang et al.,, 2023), have paved the way for more effective uncertainty quantification in probabilistic and generative models. This approach leverages conditional sampling to construct the prediction sets in the form of a union of uncertainty balls around each generated sample, thus providing a discontinuous uncertainty set. This is a fundamental step towards better coverage efficiency (small set size), as ideally, we want the prediction set to avoid uncertain low-density regions.
However, the PCP method faces two major limitations: ($i$) When the number of generated samples is small, PCP may produce overly conservative prediction sets, as it relies on the minimal distance between the true response and generated samples. ($ii$) PCP treats all generated samples equally by using equal-sized balls, which lacks the flexibility to capture complex distributional dynamics. This uniform treatment can result in inefficiencies – producing large uncertainty sets – by not fully exploiting regions where the model is more confident.
To tackle these challenges, we propose a novel Probabilistic Conformal Prediction framework by Vectorizing the non-Conformity score with Ranked samples (PCP-VCR). Specifically, our approach introduces a non-conformity score vector that ranks generated samples based on their empirical density, capturing detailed information about the error distribution across samples. By individually optimizing the quantiles for samples at the same rank, we construct uncertainty sets that are more flexible and informative. The key idea is that by leveraging the ranked structure of the non-conformity scores, our method can allocate uncertainty in a more nuanced manner. High-ranking samples from dense regions, where the model is more confident, contribute to larger prediction sets, reflecting lower uncertainty. Conversely, lower-ranking samples from sparse regions may have smaller prediction sets or be excluded altogether due to higher uncertainty. This prioritization results in tighter prediction sets that are both flexible and efficient, providing practitioners with more informative uncertainty quantification. In our numerical study, we demonstrate our method shows promising empirical performance on both synthetic and real-world datasets, compared with state-of-the-art methods. In summary, the main contributions of this paper are:
  • PCP-VCR: We present a novel method that utilizes vectorized non-conformity scores with ranked samples to enhance the efficiency of probabilistic conformal prediction.
  • Optimized Quantile Adjustment: We develop an optimization framework that individually adjusts quantile levels for each rank, leading to tighter and more informative uncertainty sets.
  • Efficient Computation: We propose an efficient heuristic algorithm to approximate the optimal quantile vector, making the method computationally practical for large datasets.
  • Theoretical and Empirical Validation: We demonstrate through theoretical analysis and empirical evaluations that PCP-VCR maintains valid coverage while significantly improving prediction set efficiency compared to baseline methods.
Related work
Other than conformal prediction, there have been several approaches proposed to quantify uncertainty such as Bayesian based methods (Blundell et al.,, 2015; Krueger et al.,, 2017; Mobiny et al.,, 2021), and ensemble based methods (Lakshminarayanan et al.,, 2017; Valdenegro-Toro,, 2019; McDermott and Wikle,, 2019; Wen et al.,, 2020). We refer to several survey papers (Gawlikowski et al.,, 2023; Abdar et al.,, 2021; He and Jiang,, 2023) for further details of works in these uncertainty quantification methods.
In terms of conformal prediction, extended works have been done focusing on CP for deterministic model. Different summary statistics of the conditional distribution have been studied, for example mean (Shafer and Vovk,, 2008; Lei et al.,, 2018) and quantiles (Romano et al.,, 2019; Kivaranovic et al.,, 2020; Alaa et al.,, 2023). In Sesia and Romano, (2021), the authors use histograms to estimate the conditional distribution and provide efficient prediction interval with approximate conditional coverage. Though the above methods all provide continuous prediction intervals, they cannot adapt to multimodality well. In Lei et al., (2013) and Lei and Wasserman, (2014), the authors first consider the aspect of density level sets and density-based clustering for providing efficient prediction set. The authors use kernel density estimator to construct the non-conformity scores and with a proper choice of bandwidth, the prediction set could be discontinuous. It is also proved that the method is asymptotically near-optimal when the number of calibration data goes to infinity. There are also some other works (Izbicki et al.,, 2019; Chernozhukov et al.,, 2021; Izbicki et al.,, 2022) that estimate the density or cumulative distribution function to construct prediction set. As a reminder, in our setting, we only need access to the random samples generated from the fitted model instead of explicit density estimation. So our setting is more generic and applicable to real applications. In Tumu et al., (2024), the authors fit shape functions (convex hull, hyperrectangle, and ellipsoid) on the calibration data and minimize the total volume of shapes to reduce the inefficiency. In Guha et al., (2024), the authors convert regression to a classification problem and can provide discontinuous prediction set with interpolation of classification scores.
Efficiency of the CP prediction set is a fundamental assessment of the quality of CP method. Some works explore the effect of different non-conformity score measures such as by scaling and reweighting the non-conformity score of each calibration data (Papadopoulos et al.,, 2002, 2011; Kivaranovic et al.,, 2020; Bellotti,, 2020; Amoukou and Brunel,, 2023). Specifically, in H. Zargarbashi et al., (2023), the authors utilize the network structure to construct a non-conformity score that incorporates neighborhood information. On the other hand, some works (Stutz et al.,, 2021; Einbinder et al.,, 2022) directly reduce inefficiency by incorporating a differentiable inefficiency measurement into the training loss. Bai et al., (2022) reformulate the conformal prediction as a constrained optimization problem. By introducing an extra parameter and enriching the search space of the optimization problem, it demonstrates better performance in efficiency. Our main idea is partially motivated by this optimization reformulation. We also gain insights from the recent development of CP for multi-target regression tasks and the definition of multivariate quantile function as discussed in (Messoudi et al.,, 2021; Sun and Yu,, 2022).
2 Preliminaries
In this section, we provide an overview of Conformal Prediction and Probabilistic Conformal Prediction (Wang et al.,, 2023), respectively, along with the key notations that will be used throughout the paper.
2.1 Conformal Prediction
Conformal Prediction (CP) is a general framework for uncertainty quantification that produces statistically valid prediction regions for any underlying point predictor, only assuming exchangeability of the data (Angelopoulos et al.,, 2023). Consider a dataset of $n$ data pairs: $\mathcal{D}=\{(X_{i},Y_{i})\}_{i=1}^{n}$, where $X_{i}\in\mathcal{X}\subseteq\mathbb{R}^{q}$ are the features and $Y_{i}\in\mathcal{Y}\subseteq\mathbb{R}^{d}$ are the responses. Given a new test pair $(X_{n+1},Y_{n+1})$, which is assumed to be exchangeable with the observations in $\mathcal{D}$, the goal of CP is to construct a prediction set $\hat{\mathcal{C}}(X_{n+1})$ based on the calibration data $\mathcal{D}$ such that, for a user-defined miscoverage rate $\alpha$, the following condition holds:
$\mathbb{P}(Y_{n+1}\in\hat{\mathcal{C}}(X_{n+1}))\geq 1-\alpha.$ (1)
This ensures that the constructed prediction set $\hat{\mathcal{C}}(\cdot)$ covers the true label with a probability of at least $1-\alpha$, where the probability is marginal over the calibration data and the test point.
To achieve this, CP relies on computing the non-conformity score, which measures the extent to which a new data point conforms to the behavior of the training data under a predictive model. For regression tasks, a typical choice of non-conformity score is the absolute residuals from a predictive model $\hat{f}$, i.e., $E_{i}=\|Y_{i}-\hat{f}(X_{i})\|$. Based on this score, the empirical quantile of the non-conformity scores is calculated, and the prediction set $\hat{\mathcal{C}}(X_{n+1})$ is constructed such that the miscoverage rate condition is satisfied.
Prediction Set Efficiency. Beyond ensuring valid coverage, efficiency is a key evaluation metric of CP’s performance, assessed by the size of the prediction set $|\hat{\mathcal{C}}(X_{n+1})|$. While a trivial prediction set that includes the entire response space would guarantee coverage, it offers little practical value for decision-making. Therefore, we aim to produce prediction sets that are as informative (i.e., small) as possible while maintaining the coverage guarantee.
2.2 Probabilistic Conformal Prediction
Probabilistic Conformal Prediction (PCP) (Wang et al.,, 2023) extends the CP framework to probabilistic (or generative) models. Given a fitted conditional probabilistic model $\hat{p}(Y|X)$, the approach involves sampling from the model rather than making a point prediction. Specifically, for each calibration point $(X_{i},Y_{i})$, we generate $K$ random samples
${\hat{Y}_{i,1},\dots,\hat{Y}_{i,K}}\sim\hat{p}(Y|X_{i}).$
The non-conformity score is then defined as the minimum distance between the true response $Y_{i}$ and the generated samples:
$E_{i}=\min_{k=1,\dots,K}\|Y_{i}-\hat{Y}_{i,k}\|.$ (2)
For a new test point $(X_{n+1},Y_{n+1})$, the generative model produces $K$ random samples ${\hat{Y}_{1},\dots,\hat{Y}_{k}}$. The prediction set is constructed as the union of $K$ regions:
$\hat{\mathcal{C}}(X_{n+1})=\cup_{k=1}^{K}\hat{\mathcal{C}}_{k},$ (3)
where each region $\hat{\mathcal{C}}_{k}$ is defined by
$\hat{\mathcal{C}}_{k}=\left\{y:\|y-\hat{Y}_{n+1,k}\|\leq Q(E_{1:n}\cup\{\infty\};\alpha)\right\}.$
Here $Q(\cdot;\alpha)$ is the $1-\alpha$ quantile of the empirical distribution of the non-conformity scores $E_{1},\dots,E_{n}$. If the error measure used is the $\ell_{2}$ norm, the resulting prediction set is a union of balls centered at each generated sample $\hat{Y}_{k}$, with a constant radius determined by the quantile $Q(\alpha)$.
A significant limitation of the PCP is its inefficiency, especially when $K$ is limited. The non-conformity score in (2) relies on the minimal distance between the true response and all generated samples, resulting in prediction sets that are unions of equal-sized balls centered at the generated samples. If $K$ is small, there’s a high chance that none of the generated samples $\{\hat{Y}_{i,k}\}_{k=1}^{K}$ are close to $Y_{i}$, leading to large nonconformity scores $E_{i}$. As a result, the quantile $Q(\alpha)$ becomes large to maintain the desired coverage level, causing the prediction sets $\hat{\mathcal{C}}(X_{n+1})$ to expand significantly and reducing their efficiency.
Figure 1: An illustrative comparison betwen PCP and proposed PCP-VCR, where PCP-VCR provides a more flexible and efficient prediction set. The non-conformity score vector $(E_{i,1},\dots,E_{i,K})^{\top}$ provides more detailed information on the underlying distribution represented by the blue area. Samples in denser regions (darker blue) are ranked higher (small $r$), indicating greater model confidence and larger $\mathcal{C}_{r}$ sizes, while low-rank scores (large $r$) capture uncertainty in sparser regions, where $\mathcal{C}_{r}$ sizes are smaller.
3 Proposed Method
In this section, we introduce a novel method for probabilistic conformal prediction, which vectorizes the non-conformity score with ranked samples, referred to as PCP-VCR.
The key idea of our method is to introduce a vectorized non-conformity score, which ranks the generated samples based on their empirical density. Unlike traditional non-conformity measures, such as the $\ell_{2}$ norm, which yield a single scalar value, our vectorized approach captures more detailed information about the distribution of errors across samples. We then jointly optimize the quantile for each ranked sample, maximizing the efficiency of the uncertainty set while preserving the validity guarantee. The uncertainty set is constructed by placing an uncertainty ball around each sample, with the radius of each ball determined by its respective optimized quantile, leading to more flexible and informative uncertainty quantification. Figure 1 illustrates a comparison between our method and the traditional PCP approach. The detailed steps of our proposed method are outlined in Algorithm 1.
Algorithm 1 PCP-VCR
1:  Input: Conditional probabilistic model $\hat{p}(Y|X)$; calibration data $\mathcal{D}$; testing feature $X_{n+1}$; coverage rate $1-\alpha$; number of random samples $K$.
2:   // Calibration
3:  Initialize $\mathcal{E}\leftarrow\varnothing$;
4:  for $(X_{i},Y_{i})\in\mathcal{D}$ do
5:     Sample $\{\hat{Y}_{i,k}\}_{k=1}^{K}\sim\hat{p}(Y|X_{i})$;
6:     $\{\hat{Y}_{i,j(r)}\}_{r=1}^{K}\leftarrow\text{Rank}~{}\{\hat{Y}_{i,k}\}_{k=1}^{K}$ by (4);
7:     $E_{i}\leftarrow$ Compute (5) for each $r$;
8:     $\mathcal{E}\leftarrow\mathcal{E}\cup E_{i}$;
9:  end for
10:  $\beta^{*}\leftarrow\operatorname*{argmin}$ of (8) calculated with $\mathcal{E}$;
11:   // Testing
12:  Sample $\{\hat{Y}_{n+1,k}\}_{k=1}^{K}\sim\hat{p}(Y|X_{n+1})$;
13:  $\{\hat{Y}_{n+1,j(r)}\}_{r=1}^{K}\leftarrow\text{Rank}~{}\{\hat{Y}_{n+1,k}\}_{k=1}^{K}$ by (4);
14:  Output: $\hat{\mathcal{C}}_{\text{VCR}}(X_{n+1};\beta^{*})$ in (6).
3.1 Vectorized Non-Conformity Score
For each calibration data $(X_{i},Y_{i})$, its vectorized non-conformity score can be constructed as follows:
  1. 1.
    Generate $K$ samples $\{\hat{Y}_{i,k}\}_{k=1}^{K}$ from $\hat{p}(\cdot|X_{i})$.
  2. 2.
    Rank these $K$ samples based on their average $m$-nearest neighbor distances, denoted by $\{\bar{D}_{i,k}\}_{k=1}^{K}$, where $\bar{D}_{i,k}$ represents the average distance from $\hat{Y}_{i,k}$ to its $m$ nearest neighbors among the generated samples:
    $\bar{D}_{i,j(1)}\leq\dots\leq\bar{D}_{i,j(K)},$ (4)
    where $j(r)$ denotes the index of the $r$-th sorted sample. Detailed computation of the average $m$-nearest neighbor distances can be found in Appendix A.
  3. 3.
    The non-conformity score vector can be obtained by
    $(E_{i,1},\dots,E_{i,K})^{\top}\in\mathbb{R}^{K},$
    where each $E_{i,r}$ is the distance between label $Y_{i}$ and the $r$-th sorted sample $\hat{Y}_{i,j(r)}$, defined as
    $E_{i,r}=\|Y_{i}-\hat{Y}_{i,j(r)}\|.$ (5)
The rationale of constructing the non-conformity score vector with ranked samples are two-fold: ($i$) By forming a vector of non-conformity scores rather than a single scalar value, we capture more detailed information about the relationship between the true label $Y_{i}$ and the generated samples. Each element $E_{i,r}$ in the vector corresponds to a specific rank $r$, reflecting different aspects of the sample distribution. This enriched information enables us to construct more flexible and informative prediction sets compared to traditional CP/PCP methods that rely on a scalar score. ($ii$) Ranking the samples based on their average $m$-nearest neighbor distances allows us to identify and prioritize samples located in denser regions of the estimated model $\hat{p}$. Samples with smaller $\bar{D}_{i,j(1)}$ are presumed to be in high-density areas and are thus ranked higher. Consequently, the non-conformity scores $E_{i,r}$ for different ranks $r$ carry varying significance: High-rank scores ($r$ is small) correspond to samples in denser regions, where the model is more confident, whereas low-rank scores ($r$ is large) represent samples in sparser regions, capturing the model’s uncertainty. By comparing non-conformity scores across different calibration data points at the same rank, we can assess discrepancies at multiple levels of the distribution. This stratified approach provides a more nuanced understanding of the model’s performance across the entire output space, leading to more accurate and adaptive uncertainty sets.
3.2 Efficiency-Enhanced Prediction Set
The prediction set can be constructed as follows: First, we compute the non-conformity score vectors $\{E_{i}\coloneqq(E_{i,1},\dots,E_{i,K})^{\top}\}_{i=1}^{n}$ for the calibration data $\mathcal{D}$. Then, we define the prediction set as follows:
$\hat{\mathcal{C}}_{\text{VCR}}(X_{n+1};\beta)=\cup_{r=1}^{K}\hat{\mathcal{C}}_{r}(\beta_{r}),$ (6)
where $\beta\coloneqq\{\beta_{r}\}_{r=1}^{K}$ and each $\hat{\mathcal{C}}_{r}(\beta_{r})$ is defined by
[…]
$Q_{r}(\{E_{1,r},\dots,E_{n,r}\}\cup\{\infty\};\beta_{r})\Big{\}},$
and $Q_{r}(\cdot;\beta_{r})$ is the $1-\beta_{r}$ quantile of the empirical distribution of the $r$-th ranked non-conformity scores across $n$ samples, $E_{1,r},\dots,E_{n,r}$. For notational simplicity, we use $Q_{r}(\beta_{r})$ throughout the paper.
Unlike PCP, which constructs its prediction set using equal-sized balls as defined in (3), the radius of each ball $\hat{\mathcal{C}}_{r}(\beta_{r})$ in (6) is rank-dependent, determined by $Q_{r}(\beta_{r})$. By introducing the quantile vector $(Q_{r}(\beta_{1}),\dots,Q_{r}(\beta_{K}))^{\top}$, we allow each rank $r$ to have its own quantile level $\beta_{r}$, providing greater flexibility in constructing the prediction set. This approach enables us to optimize the quantiles individually for each rank, leading to more efficient and tighter uncertainty sets while still ensuring valid coverage – even though the $\beta_{r}$ values are not necessarily equal to the user-defined miscoverage rate $\alpha$.
Quantile Optimization
We aim to find optimal quantile vector $\beta=(\beta_{1},\dots,\beta_{K})^{\top}$ that maximize the predictive efficiency (i.e., minimize the size of the prediction set) while ensuring the coverage validity.
To assess efficiency, we consider the total volume of all balls, which serves as an upper bound on the true size of the prediction set for computational simplicity:
$\sum_{r=1}^{K}(Q_{r}(\beta_{r}))^{d}\propto\sum_{r=1}^{K}\hat{\mathcal{C}}_{r}(\beta_{r})\geq|\hat{\mathcal{C}}_{\text{VCR}}(\beta)|,$
where $d$ is the dimensionality of the response variable $Y$. The coverage rate can be estimated by
$\frac{1}{n+1}\sum_{i=1}^{n}\hat{S}_{i}(\beta),$ (7)
where $\hat{S}_{i}(\beta)$ equals $1$ if there is at least one ball in the prediction set covers the calibration data $Y_{i}$, i.e.,
$\hat{S}_{i}(\beta)\coloneqq\max_{r=1,\dots,K}\mathbbm{1}\left\{E_{i,r}\leq Q_{r}(\beta_{r})\right\}.$
Formally, given the non-conformity score vectors, we formulate the following optimization problem:
$\min_{\beta}\left\{\sum_{r=1}^{K}(Q_{r}(\beta_{r}))^{d}\right\},~{}\text{s.t.}~{}\frac{1}{n+1}\sum_{i=1}^{n}\hat{S}_{i}(\beta)\geq 1-\alpha.$ (8)
We denote the optimal solution of this optimization problem as $\beta^{*}$. In the following theorem, we demonstrate that the prediction set $\hat{\mathcal{C}}_{\text{VCR}}(X_{n+1};\beta^{*})$ provides valid coverage:
Theorem 3.1 (Validity of PCP-VCR).
For testing data $(X_{n+1},Y_{n+1})$ that is exchangeable with calibration $\mathcal{D}$, the prediction set $\hat{\mathcal{C}}_{\text{VCR}}(X_{n+1};\beta^{*})$ provides valid coverage, i.e.,
$\mathbb{P}(Y_{n+1}\in\hat{\mathcal{C}}_{\text{VCR}}(X_{n+1};\beta^{*}))\geq 1-\alpha.$
The proof of Theorem 3.1 is presented in Appendix B.
Remark 3.2.
We note that the proof of Theorem 3.1 is non-trivial. Compared to other conformal prediction validity theorems, which typically deal with scalar non-conformity scores, the novelty here is that it extends the coverage validity to a more complex setting of vectorized and ranked non-conformity scores. For scalar non-conformity scores, the quantile is unique and can be readily determined by sorting the scalar values. However, with multivariate vectors, two primary challenges arise in computing the quantile: ($i$) The quantile is represented by a vector, in which its value is not unique ($ii$) There is no (total) ordering between vectors, meaning that vectors are not fully comparable and thus cannot be sorted as scalars are. The theorem addresses these challenges of computing valid quantiles for vectors. By defining a partial ordering between vectors and optimizing across this structure, the theorem ensures that even with this increased complexity, the prediction set maintains the desired level of validity while offering more efficient and adaptive uncertainty quantification.
Efficient Algorithm for Quantile Optimization. Solving problem (8) becomes computationally infeasible for larger values of $K$ (e.g., $K\geq 5$) due to the exponential growth of the search space. To tackle this challenge, we introduce an iterative algorithm that approximates the optimal quantile vector in linear time by simplifying the search process. The key idea is to activate one quantile at a time while setting others to default values, and then perform pairwise trade-offs between quantiles of different ranks to enhance efficiency. Specifically, we initialize the algorithm by setting $\beta_{r}=0$ for a selected rank $r$, with all other $\beta_{s}=1$ for $s\neq r$, effectively reducing the non-conformity score vector to a scalar case. We incrementally increase $\beta_{r}$ by a small step size $\epsilon$ until the empirical coverage just exceeds the target level $1-\alpha$. Subsequently, we iteratively adjust $\beta_{r}$ and other $\beta_{s}$ by trading off small increments and decrements while maintaining the coverage constraint, accepting adjustments that improve efficiency. This process has a computational complexity of $\mathcal{O}(Kn)$ (or $\mathcal{O}(K^{2}n)$ when iterating over all positions) and empirically yields an approximate solution $\hat{\beta}$ that achieves superior efficiency compared to baseline methods. The details of the algorithm can be found in Appendix C.
4 Experiments
In this section, we evaluate the empirical performance of PCP-VCR compared with several state-of-the-art baseline methods on five synthetic data, one semi-synthetic data, and four real-world datasets. We demonstrate the superior efficiency and coverage validity of PCP-VCR, especially in handling complex, multimodal, and heterogeneous distributions.
Table 1: Summary of synthetic data results.
Dataset Metric1 CQR CHR CRD R2CCP PCP PCP-VCR
Cover $0.90\pm 0.02$ $0.90\pm 0.02$ $0.89\pm 0.09$ $0.90\pm 0.02$ $0.90\pm 0.02$ $0.89\pm 0.02$
S-shape Eff $3.41\pm 0.07$ $3.43\pm 0.08$ $6.98\pm 0.79$ $3.05\pm 0.20$ $0.59\pm 0.11$ $\mathbf{0.55\pm 0.08}$
Cover $0.90\pm 0.02$ $0.91\pm 0.02$ $0.93\pm 0.03$ $0.90\pm 0.02$ $0.90\pm 0.01$ $0.88\pm 0.01$
Mix-Gaussian Eff $30.5\pm 0.00$ $31.01\pm 1.63$ $92.17\pm 8.8$ $35.09\pm 1.79$ $19.49\pm 2.44$ $\mathbf{18.43\pm 2.46}$
Cover $0.90\pm 0.01$ $0.90\pm 0.01$ $0.90\pm 0.03$ $0.89\pm 0.00$ $0.90\pm 0.01$ $0.89\pm 0.01$
Spirals Eff $20.20\pm 0.71$ $19.47\pm 0.61$ $44.48\pm 2.98$ $12.84\pm 0.31$ $3.39\pm 0.21$ $\mathbf{3.19\pm 0.18}$
Cover $0.90\pm 0.02$ $0.90\pm 0.02$ $0.90\pm 0.02$ $0.89\pm 0.00$ $0.90\pm 0.02$ $0.88\pm 0.02$
Circles Eff $1.72\pm 0.03$ $1.46\pm 0.11$ $2.91\pm 0.43$ $1.02\pm 0.01$ $0.88\pm 0.05$ $\mathbf{0.85\pm 0.05}$
Cover $0.90\pm 0.02$ $0.90\pm 0.02$ $0.90\pm 0.03$ $0.90\pm 0.02$ $0.90\pm 0.02$ $0.88\pm 0.02$
Unbalanced Cluster Eff $12.29\pm 0.17$ $12.04\pm 0.20$ $15.34\pm 1.58$ $12.83\pm 0.95$ $6.98\pm 0.64$ $\mathbf{6.79\pm 0.67}$
  • 1
    We evaluate the performance for each method with empirical coverage rate (Cover) and prediction set efficiency (Eff). Each cell contains the averaged metric value along with the standard error over 100 independent experiments.
Baselines:
We consider the following five baseline methods: ($i$) conformalized quantile regression (CQR) (Romano et al.,, 2019), ($ii$) conformal histogram regression (CHR) (Sesia and Romano,, 2021), ($iii$) probabilistic conformal prediction (PCP) (Wang et al.,, 2023), ($iv$) regression-to-classification conformal prediction (R2CCP) (Guha et al.,, 2024), and ($v$) conformal region designer (CRD) (Tumu et al.,, 2024). CQR and CHR estimate conditional quantiles and conditional histograms, respectively, to construct the prediction set. R2CCP converts regression to a classification problem and interpolates the classification scores to construct the non-conformity score for the regression problem. CRD fits shapes (hyperrectangle, ellipsoid, and convex hull) and optimizes their total volume to construct prediction sets. Among the baseline methods, CQR and CHR only produce continuous prediction sets, whereas all the other methods can provide discontinuous prediction sets. For the conditional quantile and histogram estimation of CQR and CHR, we apply the quantile regression neural network model (Taylor,, 2000) and quantile regression forest (Meinshausen,, 2006). The implementation of each baseline method is based on their public GitHub.
Implementation Configuration:
In all the following experiments, we consider the miscoverage rate $\alpha=0.1$ and evaluate the results based on the empirical coverage rate (Cover) and prediction set efficiency (Eff). The choice of hyperparameter $K$ is a trade-off between coverage efficiency and computational cost. Intuitively, with more random samples, samples can be ranked more accurately and better reflect the estimated conditional distribution, thus outputting more efficient prediction set. However, it also incurs with higher computational costs. In our experiments, we set $K=10,20,50$ for different datasets, with the specifics elaborated in the following section. In comparison, Wang et al., (2023) choose $K=1,000$ for real data analysis. We also aim to provide results of the method’s performance when the choice of $K$ is limited. We take $m=\left\lceil K/3\right\rceil$ for the $m$-nearest neighbor distance calculation. To generate conditional random samples, we implement the conditional nearest neighbor density estimation method in Izbicki and B. Lee, (2017) for $1$-dimensional synthetic data and the mixture density network (Bishop,, 1994; Rothfuss et al.,, 2019) for the other datasets as the base models. Further details of the implementation of each method, synthetic data generation, and modeling procedure can be found in Appendix D.
4.1 Synthetic Data
We evaluate the methods on five synthetic datasets: S-shape, Spirals, Circles, Unbalanced cluster for $Y\in\mathbb{R}$ and Mix-Gaussian for $Y\in\mathbb{R}^{2}$. The conditional distribution of all the datasets presents multimodality and some are also heterogeneous. We set $K=20$ for Mix-Gaussian data and $K=10$ for the other four data. Each data contains $5,000$ random observations and we split the data into $60$/$20$/$20$ for training, calibration, and testing separately.
We present the summary of the synthetic data results averaged over 100 experiments in Table 1 and highlight the most efficient method for each data. Overall, all methods achieve valid coverage rates, while PCP-VCR produces the most efficient prediction set. The visualizations of the prediction sets of different methods on these data are presented in Appendix E. We can observe that PCP and PCP-VCR can consistently provide discontinuous prediction sets and, thus are much more efficient than other methods for multimodal settings. Compared with PCP, PCP-VCR provides a more flexible and locally-adaptive prediction set, i.e., the coverage region of PCP-VCR is larger in the high-density region and smaller in the low-density region, especially in the Mix-Gaussian setting. This is the benefit that vectorized non-conformity scores provide compared with a scalar.
Figure 2: Comparison of MNIST images sampled from the prediction set provided by PCP-VCR and PCP, respectively. Each row corresponds to different values of $X$. We observe that samples of PCP-VCR align more closely with the target $X$ compared to PCP. It indicates that PCP-VCR provides more efficient prediction sets with fewer wrong images covered.
4.2 MNIST: Digit Image Generation
We design a semi-synthetic experiment with the MNIST image dataset (Deng,, 2012) to assess the applicability and interpretability of PCP-VCR to high-dimensional data. Conformal prediction in high-dimensional data, such as images, texts, and audio, presents a unique challenge for interpreting prediction sets, as it is inapplicable to iterate over the entire space of target variables.
Consider the setting that given $X\in\{0,\dots,9\}$, the corresponding Y is an image of digit $X$. We first train a variational-autoencoder (VAE) model and create the prediction set in the embedding space. For the conditional sampling process, we assume we have access to an independent sampler. For any given $X$, it will output random images of digit $X$ that are not presented in the training/calibration/testing process to ensure independence. We then embed the sampled images with the trained encoder and construct the prediction set in the embedding space. Ahead of training, we split part of the data for the sampler so it is independent of the following process. The reason that we design this sampler instead of training a conditional VAE (cVAE) is that the embedding space of VAE tends to be more clustered than cVAE, thus the prediction set can be better demonstrated for illustrative purposes.
Figure 2 shows samples from the prediction set provided by PCP and PCP-VCR with $K=40$ conditioning on different $X$ values in each row. For each $X$, we randomly sample 9 images from the prediction set of PCP and PCP-VCR. Both of the methods achieve valid coverage rates. We can observe that the sampled images of PCP-VCR align closer with the target image for each $X$, indicating a tighter and more efficient prediction set of PCP-VCR. Quantitatively, as it is not applicable to directly measure the size of the prediction set, we propose to evaluate the miscoverage rate. It measures the empirical probability of an incorrect image falling into the prediction set. Under $K=40$, PCP-VCR attains an average miscoverage rate of $0.16\pm 0.02$ (standard error) over 30 experiments, compared to $0.22\pm 0.03$ for PCP. In Figure 3, we present the comparison of the miscoverage rate of PCP and PCP-VCR under different choices of $K$. It demonstrates that PCP-VCR has a lower miscoverage rate than PCP consistently. This also explains why the random images of PCP-VCR aligns more closely with the target as shown in Figure 2
Figure 3: Comparison of miscoverage rate of PCP and PCP-VCR for MNIST data under different $K$. The blue and red bars represent PCP and PCP-VCR results, respectively, with black lines showing the standard error over 30 experiments.
Table 2: Summary of real data results.
Dataset Metric 1 CQR CHR CRD R22CCP PCP PCP-VCR
Cover $0.97\pm 0.00$ $0.90\pm 0.02$ $0.85\pm 0.16$ $0.33\pm 0.00$ $0.93\pm 0.01$ $0.89\pm 0.02$
Fb1 Eff $19.78\pm 1.95$ $9.50\pm 0.81$ $58.31\pm 62.27$ $7.96\pm 0.00$ $9.05\pm 9.01$ $\mathbf{4.84\pm 2.38}$
Cover $0.97\pm 0.01$ $0.90\pm 0.01$ $0.90\pm 0.02$ $0.52\pm 0.25$ $0.93\pm 0.01$ $0.89\pm 0.01$
Fb2 Eff $19.34\pm 1.43$ $9.52\pm 0.90$ $44.16\pm 33.98$ $11.04\pm 2.00$ $8.48\pm 8.83$ $\mathbf{4.35\pm 1.90}$
Cover $0.97\pm 0.01$ $0.90\pm 0.01$ $0.89\pm 0.01$ $0.32\pm 0.58$ $0.92\pm 0.02$ $0.90\pm 0.01$
Blog Eff $22.01\pm 2.01$ $9.71\pm 1.21$ $24.56\pm 8.45$ $11.04\pm 2.00$ $11.04\pm 2.50$ $\mathbf{3.96\pm 2.23}$
Cover $0.91\pm 0.01$ $0.90\pm 0.01$ $0.90\pm 0.01$ $0.90\pm 0.02$ $0.90\pm 0.01$ $0.89\pm 0.01$
Bike Eff ($10^{-3}$) $1510\pm 10$ $4.7\pm 0.10$ $10.71\pm 3.78$ $0.79\pm 0.09$ $0.73\pm 0.26$ $\mathbf{0.57\pm 0.06}$
  • 1
    Each cell contains the averaged metric value along with the standard error over 30 independent experiments.
(a) PCP-VCR
(b) PCP
(c) CHR
(d) R2CCP
Figure 4: Comparison of prediction sets across four methods that have the highest efficiency on Bike dataset. The colored area indicates the prediction set, and the red marker is the target location.
4.3 Real Data
We apply PCP-VCR along with baseline methods on four real-world datasets: blog feedback (Blog) Buza, (2014), facebook comment volumne Singh, (2015) variant 1 (Fb1) and variant 2 (Fb2), and New York City bike sharing (Bike). For each dataset, we randomly sample $30,000$ samples for training, $1,000$ samples for calibration and $1,000$ for testing. We set $K=50$ for the Bike dataset, and $K=20$ for the rest datasets. The summary of the real data results averaging over 30 experiments is presented in Table 2. Overall, PCP-VCR still performs the best, and it achieves the best efficiency while maintaining the coverage rate. It is noted that for the first three datasets, R2CCP suffers from the undercover issue and has a large coverage rate variance across experiments.
We present a visual demonstration of the prediction set generated by our method, PCP, CHR and R2CCP in Figure 4, respectively. The blue region indicates the prediction set of each method and the red star marker indicates the target location. Both PCP and our method deliver discontinuous prediction regions, whereas CHR and R2CCP provide rectangle-shaped continuous sest. We also notice that, compared with PCP, PCP-VCR assigns larger circles’ radius to high-density regions while nearly vanishing the circles’ radius in low-density regions. Specifically, the circle at the top of the map lies in a low-density region, it almost shrinks to a point for PCP-VCR, whereas it remains large for PCP. Thus, PCP-VCR leads to significant improvements in flexibility and efficiency of prediction set.
5 Discussion and Conclusion
In this paper, we proposed PCP-VCR, a novel conformal prediction method for conditional generative models utilizing vectorized non-conformity scores with ranked samples to enhance the efficiency of the prediction set. By ranking the conditional samples, our method explores and captures the structure among the random samples and incorporates it into non-conformity score vector. By separately optimizing each entry of the non-conformity score vector, it enables the creation of a flexible and data-adaptive prediction set. Specifically, PCP-VCR outputs discontinuous prediction set in the form of a union of regions centered at random samples, where the radius of each region is adapted to the model’s uncertainty level of each sample. Additionally, we proposed a computationally efficient algorithm for approximating the optimal quantile vector, ensuring the method practical even for large datasets.
Through extensive experiments on both synthetic and real-world datasets, PCP-VCR consistently demonstrated superior coverage efficiency while maintaining coverage validity. In the MNIST experiment, we showcased the potential power of PCP-VCR for conformal prediction in general generative model settings, illustrating its ability to improve the reliability of predictions in modern high-stakes applications.
Appendix A Average $m$-Nearest Neighbor Distance
The average $m$-nearest neighbor distance for a generated sample $\hat{Y}_{i,k}$ can be computed as follows:
  1. 1.
    (Optional) Sample a sequence of reference samples $\{\tilde{Y}_{i,j}\}_{l=1}^{M}\sim\hat{p}(\cdot|X_{i})$ for ranking estimation. Otherwise, set $\tilde{Y}_{i,l}=\hat{Y}_{i,l}$ for $l=1,\dots,K$ and $M=K$.
  2. 2.
    Compute the pairwise distances of generated samples among reference samples:
    $D_{i}(k,l)\coloneqq\|\hat{Y}_{i,k}-\tilde{Y}_{i,l}\|,~{}\text{for}~{}k\in\{1,\dots,K\},\ l\in\{1,\dots,M\}.$
  3. 3.
    Sort the distances to $\hat{Y}_{i,k}$ such that:
    $D_{i}(k,(1))\leq\dots\leq D_{i}(k,(M)),$
    where $D_{i}(k,(l))$ denotes the $l$-th smallest distance to $\hat{Y}_{i,k}$.
  4. 4.
    Calculate average $m$-nearest neighbor distance:
    $\bar{D}_{i,k}=\frac{1}{m}\sum_{l=1}^{m}D_{i}(k,(l)).$
Appendix B Proof of Theorem 3.1
To prove the theorem, we first need to define the partial ordering between vectors, then we will define the empirical multivariate quantile function w.r.t. this partial ordering.
Definition B.1 (Vector partial order).
We define a partial order $\succ$ for $K$-dimensional vectors $u,v$ as the following,
$u\succ v\quad\textup{i.f.f.}\quad\forall r\in\{1,\dots,K\},u_{r}>v_{r}.$ (9)
Proof.
Based on the definition, we notice that
$\mathbbm{1}\{u\succ v\}=1-\max_{r=1,\dots,K}\mathbbm{1}\{u_{r}\leq v_{r}\}.$
We can rewrite the constraint in Equation 8 as
$\begin{split}\frac{1}{n+1}\sum_{i=1}^{n}\hat{S}_{i}(\beta)&=\frac{1}{n+1}\sum_{i=1}^{n}\max_{r=1,\dots,K}\mathbbm{1}\left\{E_{i,r}\leq Q_{r}(\beta_{r})\right\}\\ &=\frac{1}{n+1}\sum_{i=1}^{n}(1-\mathbbm{1}\{E_{i}\succ Q(\beta)\})\\ &\geq 1-\alpha,\end{split}$ (10)
where we let $Q(\beta)=(Q_{1}(\beta_{1}),\dots,Q_{K}(\beta_{K}))^{T}$ denote the vector of quantiles for simplicity. We can then observe the following inequality based on the last inequality in Equation 10,
$\begin{split}\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\{E_{i}\succ Q(\beta)\}<\alpha\end{split}$ (11)
We can then define the empirical multivariate quantile function as
$\hat{Q}(\{E_{1},\dots,E_{n}\},1-\alpha)$ $=\operatorname*{argmin}_{\beta\in\mathbb{R}^{K}}\sum_{r=1}^{K}(Q_{r}(\beta_{r}))^{d}$ (12)
s.t. $\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\{E_{i}\succ Q(\beta)\}<\alpha$ (13)
We let $\beta^{*}=\hat{Q}(\{E_{1},\dots,E_{n}\},1-\alpha)$, then for a testing data pair $(X_{n+1},Y_{n+1})$ that is exchangeable with $\mathcal{D}$, we have $\mathbb{P}(E_{n+1}\succ Q(\beta^{*}))<\alpha,$ by definition of $Q(\beta^{*})$. We observe that its complement set achieves the target coverage, i.e,
($\mathbb{P}(\{E_{n+1}\succ Q(\beta^{*})\}^{c})=\mathbb{P}(\cup_{r=1}^{K}\{E_{n+1,r}\leq Q_{r}(\beta^{*}_{r})\})\geq 1-\alpha.$
By the definition of $\hat{\mathcal{C}}_{\text{VCR}}(X_{n+1};\beta)$, we have the following equivalence,
$\cup_{r=1}^{K}\{E_{n+1,r}\leq Q_{r}(\beta^{*}_{r})\}\Longleftrightarrow\{Y_{n+1}\in\hat{\mathcal{C}}_{\text{VCR}}(X_{n+1};\beta^{*})\}.$
Thus we prove that $\mathbb{P}(Y_{n+1}\in\hat{\mathcal{C}}_{\text{VCR}}(X_{n+1};\beta^{*}))\geq 1-\alpha$. ∎
Appendix C Efficient Algorithm for Quantile Optimization
Deriving $\beta^{*}$ for small $K$ and $n$ can be done by grid search. However, grid search is no longer applicable even for $K\geq 5$. In this section, we propose an empirical method that computationally efficiently approximates $\beta^{*}$ with linear complexity $\mathcal{O}(Kn)$ ($\mathcal{O}(K^{2}n)$ if apply the algorithm on all random sample positions).
The key idea of the approximation method is motivated by the “trade-off” between quantiles for different ranks. Intuitively, if one quantile is decreased then another quantile has to be increased to maintain coverage validity. During this trade-off, the efficiency changes accordingly. We present a simple demonstration of this idea for $K=2$ in Figure 5. The colored region represents the feasible region in 8 with darker color indicating higher efficiency, and the red point shows the position of $\beta^{*}$. We can observe ($i$) a step-wise pattern on the boundary of the feasible region and ($ii$) $\beta^{*}$ appears on the boundary. The step-wise pattern corresponds to the trade-off between $\beta_{1}$ and $\beta_{2}$, which leads to different efficiency. Additionally, the feasible region is in a convex shape, also indicating the trade-off relationship between quantiles. Thus to approximate the optimal solution, we only need to find the boundary of the feasible region, and iteratively do the trade-offs between different quantiles.
Figure 5: Feasible region and optimal solution of optimization problem in Equation 8 for $K=2$. The colored region shows the feasible region of $\beta$ with darker color indicating higher efficiency and the red point shows the position of $\beta^{*}$.
To be more specific, the algorithm initializes a $\beta_{k}=0$ and sets all the other entries of $\beta$ to be 1. Then, we gradually increase $\beta_{k}$ until the empirical coverage is just achieved above the target coverage. The intuition behind such initialization is that it reduces the non-conformity score vector to the simple scalar case, as observed in the following equation,
$\hat{S}_{i}(\beta)\coloneqq\max_{r=1,\dots,K}\mathbbm{1}\left\{E_{i,r}\leq Q_{r}(\beta_{r})\right\}=\mathbbm{1}\left\{E_{i,k}\leq Q_{k}(\beta_{k})\right\}$
for $\beta_{k}\neq 1$ and $\beta_{j}=1$ for all $j\neq k$. Thus there exists such $\beta_{k}$ that $\beta$ can still satisfy empirical coverage with only $\beta_{k}\neq 1$. Let $\epsilon>0$ denote a small increment, we then randomly pick another entry $\beta_{j}$ for $j\neq k$ and do the trade-off between $\beta_{j}$ and $\beta_{k}$. We set $\beta_{k}=\beta_{k}+\epsilon$, and decrease $\beta_{j}$ until the empirical coverage is satisfied. Elementally, this process mimics the step-wise pattern, where we increase $\beta_{k}$ and decrease $\beta_{j}$ to do the trade-off and examine if the efficiency is improved. We iterate this process for different $j$ until reach the budget. We show that empirically, the approximated solution $\hat{\beta}$ still achieves superior efficiency than baseline methods. We refer to Algorithm 2 for the complete details of the empirical approximation method. As a remark for practice, we set $\epsilon=1/n$ as default. But in cases, if the quantiles are not updated after several iterations, one may consider increasing $\epsilon$ for a larger trade-off.
Algorithm 2 Efficient Optimial Quantiles Search
1:  Input: Non-conformity scores $\{E_{i}\}_{i=1}^{n}$; initial position $k$; budget $B$; increment $\epsilon$.
2:  Initialize:
$\beta=\{\beta_{1},\dots,\beta_{K}\}\leftarrow\{\underbrace{1,1,\dots,1}_{\text{k-1}},0,\underbrace{1,\dots,1}_{\text{K-k}}\}.$
3:  while $\frac{1}{n+1}\sum_{i=1}^{n}\hat{S}_{i}(\beta)>1-\alpha$ do
4:     $\beta_{k}\leftarrow\beta_{k}+\epsilon$;
5:  end while
6:   // Conducting trade-offs
7:  for iter $\leftarrow 1$ to $B$ do
8:     sample $j\in\{1,\dots,K\}$ s.t. $j\neq k$;
9:     $\beta_{k}^{\prime}\leftarrow\beta_{k}+\epsilon$, $\beta_{j}^{\prime}\leftarrow\beta_{j}$;
10:     Set $\beta_{k}^{\prime}$, $\beta_{j}^{\prime}$ as the $k$-th and $j$-th entry of $\beta$;
11:     while $\frac{1}{n+1}\sum_{i=1}^{n}\hat{S}_{i}(\beta)<1-\alpha$ do
12:        $\beta_{j}^{\prime}\leftarrow\beta_{j}^{\prime}-\epsilon$;
13:     end while
14:      // Check if efficiency is improved
15:     $r\leftarrow(Q_{k}(\beta_{k}))^{d}+(Q_{j}(\beta_{j}))^{d}$, $r^{\prime}\leftarrow(Q_{k}(\beta_{k}^{\prime}))^{d}+(Q_{j}(\beta_{j}^{\prime}))^{d}$;
16:     if $r>r^{\prime}$ then
17:        continue
18:     else
19:        Set $k$-th and $j$-th entry of $\beta$ to $\beta_{k}$ and $\beta_{j}$;
20:     end if
21:  end for
22:  $\hat{\beta}\leftarrow\beta$;
23:  Output: Approximated solution $\hat{\beta}$.
Figure 6: Comparison of prediction set ($\alpha=0.1$) generated by different methods on $1$-dimensional synthetic data: S-shape, Circles, Spirals, and Unbalanced clusters (from top to bottom row). Blues lines: the predictive intervals of each method; Black dots: test points that are covered by the predictive sets; Reds dots: test points not covered. We plot the figure with $100$, $500$, $400$, and $500$ random samples for each data, respectively.
Figure 7: Comparison of prediction set ($\alpha=0.1$) generated by different methods on $2$-dimensional synthetic mixture Gaussian data. The mixture Gaussian data is generated with 2 components with weights 0.7 and 0.3 from left to right, respectively. The blue shade indicates the true conditional density and the rectangles and circles denote the prediction sets of different methods. In the left subplot, we present the rectangle-shaped prediction set generated by CQR, CHR and R2CCP. In the right subplot, we present the circle-shaped prediction regions provided by PCP-VCR and PCP.
Figure 8: Ratio of PCP prediction set length over PCP-VCR prediction set length under different $K=3,5,10,15,20$ and number of calibration data for S-shape data. The blue curve represents the mean, and the shaded area indicates the standard error of the ratio. We observe that the ratio remains consistently above 1, demonstrating that PCP-VCR consistently achieves better efficiency.
Appendix D Implementation Details
For the estimation of conditional distribution for CQR and CHR, we implemented two base models, including quantile regression neural network (QNN) and quantile regression forest (QRF). We used QNN as the base model for CHR and CQR in the synthetic data experiments and QRF for the real data experiments. For the conditional probabilistic models for PCP and PCP-VCR, we conduct nearest neighbor regression (NN) for $1$-dimensional synthetic data and mixture density network (MDN) for $2$-dimensional synthetic data. For the real data experiments, we use QRF as the base models for PCP and PCP-VCR, the same with CHR and CQR.
  • Quantile regression neural network (QNN). The network comprises three fully connected layers with ReLu activation functions. We used pinball loss to estimate the conditional quantiles for a grid of quantiles ranging from $0.01$ to $1.0$ with step size $0.01$. The dropout rate is set to be $0.2$. We train the model with 100 epochs using Adam optimizer with a learning rate equal to $0.001$.
  • Quantile regression forest (QRF). We set the minimum number of samples required to split an internal node to $50$ and set the total number of trees to $100$.
  • Nearest neighbor regression (NN). We conduct NN with Python package flexcode and set the number of the nearest neighbor as $50$ with a maximum of $31$ basis functions.
  • Mixture density network (MDN). The implementation of MDN is conducted with Python package cde. We set the number of Gaussian mixture components as $10$ and the hidden size as $(100,100)$. The other parameters are set to the default. We train the MDN model for $1,000$ epochs with a dropout rate of $0.1$.
Appendix E Experiments Results
In this section, we present the additional experimental results for synthetic data. We present the visualizations of the prediction set provided by different baseline methods on S-shape, Circles, Spirals, Unbalanced clusters, and mixture-Gaussian datasets in Figure 6 and 7. For mixture-Gaussian data, we generate $X\sim\textup{Uniform}(0,1)$, and let $Y|X\sim 0.7\textup{N}((X,0)^{T},I)+0.3\textup{N}((5+X,0)^{T},I)$ for identity matrix $I\in\mathbb{R}^{2\times 2}$. In the right subplot of Figure 7, we present the comparison of the prediction set of PCP and PCP-VCR. A key observation here is that in the low-density region (right component), PCP-VCR has a smaller radius for almost every circle, whereas in the high-density region (left component), PCP-VCR has a larger region covered. This demonstrates the ability of PCP-VCR to adapt to the local density by leveraging ranked samples.
Here, we also investigate the effect of $K$ on coverage efficiency. We provide an example in Figure 8, which presents the ratio of PCP prediction set length over PCP-VCR prediction set length for S-shape data under different choices of $K=3,5,10,15,20$ and different numbers of calibration data. The curve plot represents the mean, and the shaded area indicates the standard error of the ratio over 100 experiments. Under all settings, the ratio consistently remains above 1, showing that PCP-VCR provides a more efficient prediction set. Additionally, the ratio decreases as $K$ increases, suggesting that the benefit of PCP-VCR decreases with larger $K$.